OpenStudy (anonymous):

A ball rolls down an inclined plane such that the distance (in centimeters) that it rolls in t seconds is given by s(t) = 2t^3 + 3t^2 + 4 for 0<=t<=3 At what time is the velocity 30cm/s?

OpenStudy (anonymous):

I've got the first part figured out (velocity at t = 2), but using the same equation I used for the previous problem doesn't work, dimbeni.

OpenStudy (anonymous):

v(t)=6t^2+6t 2t^3+3t^2+4=30 -> 2t^3+3t^2-26=0

OpenStudy (anonymous):

Yup. I got (-1 + sqrt(24)) / 2 as the answer but it doesn't seem right.

OpenStudy (anonymous):

Ok, sorry. I was mistaken. It's v(t)=30. Then t^2+t-5=0 I get (-1+sqrt(21))/2

OpenStudy (anonymous):

So I'm right? Sweet...!

OpenStudy (anonymous):

Good answer yet again! Thanks dimbeni...:D